OpenStudy (anonymous):

HELP PLEASE x + y = k 2x + 3y = k + 1 The lines intersect at 3x + 2y = A 5x + y = B ax + by = c dx + ey = f

OpenStudy (e.cociuba):

Solve the equations for the solutions, x and y. x + y = k 2x + 3y = k + 1 The lines intersect at (?,?) ----------------------------------- Multiply thru the 1st equation by 2: 2x + 2y = 2k 2x + 3y = k+1 --------------- Subtract the 1st from the 2nd, then solve for "y": y = 1-k ------- Substitute into x+y = k to solve for "x": x + 1-k = k x = 2k+1 ---------------------- Point of intersection: (2k+1 , 1-k)

OpenStudy (anonymous):

I tried that, it's wrong.

OpenStudy (anonymous):

BTW these are three different problems.

OpenStudy (mertsj):

x= 2k-1

OpenStudy (mertsj):



OpenStudy (mertsj):

(2k-1,1-k)

OpenStudy (anonymous):

Minus? Let me plug it in.

OpenStudy (anonymous):

Cool, thanks

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Could you help with the other two? Please

OpenStudy (mertsj):

1. Multiply the second equation by -2 2. Add the result to the first equation 3. Solve for x